OpenStudy (anonymous):

Solve the equation for m: A = 2(m – w). A. m = (A + w/2) B. m = A + w/2 C. m = 2(A + w) D. m = A/2 + w

OpenStudy (anonymous):

A/2 = m - w Divide by 2 A/2 + w = m so, D

OpenStudy (anonymous):

Thaaanks :)